A 0.23?F capacitor is connected across an AC generator that produces a peak voltage of 9.00V .

A) What is the peak current through the capacitor if the emf frequency is 100 Hz?

B) What is the peak current through the capacitor if the emf frequency is 100 kHz?
Answer #1

A)

the capacitive reactance is

XC = 1/ 2? * f * C
= 1/(2? * 100 * 0.23 x 10^-6) = 6923.289 ohm
Imax = Vmax/Xc

            = 9.0 V/6923.289 ohm

           = 1.29x 10^-3 A = 1.29 mA

B) XC = 1/ 2? * f * C
= 1/(2? * 100 x 10^3 * 0.23 x 10^-6) = 6.92ohms
Imax = Vmax/Xc

       = 9.0 V/ 6.92 ohm

        = 1.3 A